# ITCMA: A Generative Agent Based on a Computational Consciousness Structure

### v1

@@ Abstract.
Large Language Models (LLMs) still face challenges in tasks requiring understanding implicit instructions and applying common-sense knowledge. In such scenarios, LLMs may require multiple attempts to achieve human-level performance, potentially leading to inaccurate responses or inferences in practical environments, affecting their long-term consistency and behavior. This paper introduces the Internal Time-Consciousness Machine (ITCM), a computational consciousness structure. We further propose the ITCM-based Agent (ITCMA), which supports behavior generation and reasoning in open-world settings. ITCMA enhances LLMs’ ability to understand implicit instructions and apply common-sense knowledge by considering agents’ interaction and reasoning with the environment. Evaluations in the Alfworld environment show that trained ITCMA outperforms the state-of-the-art (SOTA) by 9% on the seen set. Even untrained ITCMA achieves a 96% task completion rate on the seen set, 5% higher than SOTA, indicating its superiority over traditional intelligent agents in utility and generalization. In real-world tasks with quadruped robots, the untrained ITCMA achieves an 85% task completion rate, which is close to its performance in the unseen set, demonstrating its comparable utility in real-world settings.
@@ 1. Introduction
The aim of researchers working with generative agents is to provide a framework for behavior in open worlds: a framework in which agents can interact with other agents and respond to changes in the environment. Generative agents take the current environment and past experiences as input and generate behavior as output. The foundation of such behavior is a novel agent architecture that combines LLMs with mechanisms for synthesizing and retrieving relevant information, thus providing conditions for the output of language models. Without these mechanisms, LLMs could still output behavior for agents, but the agents might not react correctly based on past experiences, fail to make crucial inferences, or struggle to maintain long-term consistency [3].
@@
As the interaction cycles of LLM-based agents grow, two issues arise with this structure. The first issue relates to the length of the historical record. LLM-based agents process previous interactions in natural language format, appending the historical record to each subsequent input. As these records extend, they might exceed the constraints of the Transformer architectures most LLM-based agents rely on. In such cases, the system might truncate certain content. The second issue is the difficulty of extracting relevant memories. As agents accumulate a large volume of historical observations and action sequences, their memory burden increases continuously. This makes establishing connections between relevant topics increasingly challenging, potentially resulting in the agent’s responses being inconsistent with the ongoing context [4].
Due to the current limitations of LLMs acting as the ”brain,” most research on LLM-based agents focuses on augmenting them with an additional layer of structures, such as LangChain [5], to optimize them. However, there is still considerable debate on how to establish such a structure.
Before the LLM, research approaches to artificial general intelligence are from the perspectives of connectionism, behaviorism, and symbolism. Connectionism focuses on the study of neural networks, arguing that artificial general intelligence will eventually be built up from a large number of neural networks; symbolism focuses on the comparison between the human brain and computer, arguing that cognition is eventually something that can be computed; and behaviorism regards intelligence as a stimulus-response model, and believes intelligence is a type of reaction machine [6; 7]. In fact, these approaches share a common research orientation: the simulation of human intelligence. In philosophy, this is often described as a simulation of the structure of human consciousness.
The most typical work here comes from cognitive science and philosophy of mind, a branch of them pays special attention to the problem of consciousness. Researchers in it discuss the problem of consciousness and human intelligence and often draw inspiration from phenomenology, especially Merleau-Ponty’s phenomenology of perception [8; 9; 10]. ”Intelligence” from Merleau-Ponty’s perspective cannot be considered as something independent from a human being, nor as an automaton detached from the world, nor can it be simply linked to some action, some body mechanism, or some specific consciousness [11]. Thompson further developed Merleau-Ponty’s theory. He believed that mental life consists of three persistent and intertwined physical activity modes: self-regulation, sensorimotor coupled with the world, and interaction between subjects [12]. Self-regulation is essential for sentience. It is obvious in emotions, feelings, and some states, such as waking or sleeping. Sensorimotor coupled with the world is expressed in perceptions, emotions, and actions. Interaction between subjects is the cognition and emotional experience of the self and others [9].
Implicit in all of this is the presupposition that intelligence cannot exist without consciousness. That is, intelligence is not the stimulus-response mode claimed by behaviorism (represented by reinforcement learning in AI research), nor is it an “emergence from neuroelectricity” claimed by connectionism (represented by deep learning in artificial intelligence research). An expression of mechanical determinism is not applicable in the study of intelligent structures. To design a more effective generative agent structure, it is necessary to consider an imitative computational consciousness structure that is not based on the underlying principles at the present time when the basic structure of consciousness is not clear. That is, to realize the algorithm simulation of intelligence, a computable structure that can be equivalent to consciousness to a certain extent first needs to be proposed.
The main focus of this paper is to optimize the mechanism by which generative agents generate behavioral outputs based on their environment and experiences. Our main contributions are as follows:
@@
  • To enhance the foundational structure of generative agents, making them more intelligent and interpretable, we propose the Internal Time-Consciousness Machine (ITCM) as a reconstruction of the underlying architecture of generative agents. This computational consciousness structure can assist generative agents in handling complex tasks with greater flexibility and intelligence while enhancing their interpretability, making their behaviors easier to understand and predict.
@@
  • Technically, based on the proposed ITCM structure, we introduce an ITCM-based agent (ITCMA) and validate its effectiveness in both of the life scenario environments described by text, as well as in real-world robotic environments.
@@ 2.1. LLMs as agents
The emergence of LLMs has brought new avenues for intelligent agents. In recent years, there has been a proliferation of LLM-based agent architectures [13; 14; 4]. These architectures primarily focus on two aspects: planning and tools [15].
@@ 2.2.1. Integrated Information Theory
The central claim of integrated information theory (IIT) is that a physical system is conscious when and only when it is the maximum of the integrated information $\phi$. As defined by the IIT, integrated information can be roughly described as a measure of the extent to which a system is causally bound to its own past and future states and the dependence of these constraints on the causal interconnectedness between system components. If a system has more integrated information than any overlapping system (e.g., a smaller system that is a part of it or any larger system containing it), then this system has the maximum comprehensive information. Consciousness needs the largest $\phi$, not just a non-zero value of $\phi$ (this is called the exclusion theorem). The IIT also claims that consciousness is identical to the maximum $\phi$ [26; 27].
The IIT has gained a large number of followers. However, it has also been met with some philosophical objections [28]. One of these objections is that consciousness intuitively looks like an intrinsic property, but maximal $\phi$ is an extrinsic property. Therefore, if this intuition is correct, then consciousness cannot be the same as maximal $\phi$ [29].
@@ 2.2.2. Consciousness Turing Machine
According to Blum [31], the basis of consciousness in the CTM uses Baar’s theater of consciousness hypothesis, which likens consciousness to the performance of dramatic actors on a stage of working memory. Their performance occurs under the observation of a group of spectators sitting in the dark (unconscious processors). The stage of the GWT is represented by short-term memory (STM) that contains the content of the CTM consciousness at any moment. The spectators are represented by powerful processors. Each processor has its own expertise, and these processors make up the CTM’s long-term memory (LTM). These processors make predictions and receive feedback from the CTM’s world. The learning algorithms inside each processor improve the processor’s behavior based on this feedback.
3. The Internal Time-Consciousness Machine (ITCM)
@@
It is undeniable that the IIT and CTM suggest ways to build algorithms based on different aspects of consciousness research. However, the question remains as to whether there is truly a computation-based theory of consciousness, even though there are many similarities between the brain and the structure of the digital computer. Thagard [32] summarizes seven challenges to the study of representation computation: (1) the brain is not “computational” in the standard sense of the term; (2) rationality is not completely independent of emotion; (3) the study of consciousness; (4) embodied cognition; (5) whether the unit of analysis of cognition should include the environment; (6) the dynamical system of the mind; and (7) the social nature of intelligence. Notably, the asymmetry between cognition and computation occupies the first position in this list of challenges.
In cognitive science, perception proceeds through three basic forms of development: sensation, perception, and representation. Sensation is the reflection of the individual properties and characteristics of objects in the human brain (with the aid of the body’s nervous system and sensory organs). Perception is the synthesis of various sensations and the reflection of objects as a whole in the human brain. On the basis of perception, a representation is produced. These representations are reproduced through recollection and association and are formed on the basis of multiple perceptions of the same thing or similar things. Representation is the process by which an image of an object that has been perceived is reproduced mentally when the object is not present in front of the individual [33]. In fact, Merleau-Ponty’s [11] study of the perceptual field falls conceptually in the middle of these three forms, and there is a striking similarity between sensation and the elements in the perceptual field, and between representation and the content of the phenomenal field above the perceptual field. This certainly provides some degree of confirmation of Merleau-Ponty’s view of the coherence of consciousness and perception.
Does this mean that if we focus only on constructing a model of pseudo-consciousness that can correspond to intelligence, we can turn our attention to first-person phenomenology? We discuss this question in more detail in Appendix A.
3.1. Basic Structure of the ITCM
3.1.1. Perceptual and Phenomenal Fields
A “field” is a functional and exclusive activity location of conscious experience, as identified by Husserl [34]. Merleau-Ponty [11] collectively referred to the various experiential “fields” established by Husserl as “phenomenal fields.” Merleau-Ponty argues that consciousness works through one part of the phenomenal field, namely, the perceptual field. Only in the perceptual field can silent experience be brought into meaningful expression, completing the actual presentation of consciousness and the conscious activity of the final representation. According to Merleau-Ponty, the perceptual subject is always in the midst of other objects; it is always a part of the field. Each specific experience can only be unified through its own field. Without a mechanism such as the perception field, our perception would not be continuous at all.
What is the content of the perceptual field? According to Husserl’s [34] theory of consciousness, the stream of monads is the most basic unit of meaning in the phenomenal field; a continuous stream of monads forms our perceptions. Libet [35] provides the strongest explanation for this: “We have shown experimentally that not all nerve cell activities give rise to a conscious experience. For example, a short train of stimulus pulses to sensory cortex elicits responses of many nerve cells without any subjective experience” (p. 166). A single cell or cell group cannot be the birthplace of conscious experience. Only when a monad becomes part of a monad stream (demonstrating continuous extensibility) can consciousness be constructed, activated, and displayed.
Therefore, let us imagine a spherical coordinate system: the perceptual subject (that is, the agent) is always in the middle of the sphere and facing forward, and the object that can be perceived is described by the set of spherical coordinates ($\theta$, $\varphi$, $\gamma$). In a single static field, such a perceived object is a monad.
When we obtain these sets of represented monads, the spherical coordinate system that can be transformed by the agent’s actions becomes the agent’s phenomenal field. Considering some simulation requirements, we can further refine this rough model. In Merleau-Ponty’s concept, the phenomenal field is a larger concept than the perceptual field. The perceptual field is influenced by the material being perceived, and perception is further influenced by sensation.
Visual examples can be used to illustrate some concepts. It is obvious that infants do not reflect on what they see. They take what they see and consider it real. However, an infant’s range of consciousness is not limited to visual stimuli. For objects that are not within their perceptual range, infants with experience will understand the possible changing conditions of these objects [36]. For example, if we often use some “sleight of hand” tricks to change a piece of candy from the left hand to the right, the infants will eventually go directly to the right hand to obtain the candy.
This implies two situations. The renewal of perception is mandatory in non-reflective consciousness; that is, for the subject, this is “fact.” In addition, before perception is forced to update, there will still be an impression of the object in the phenomenal field, and it is not static. It will move and transform according to the original experience until perception forces an update.
We can make some changes to the proposed definitions for a computer simulation: the phenomenal field is a spherical coordinate system centered on the subject, which contains many monads, and it can make some predictive changes to these monads according to experience; the perceptual field is a part of the phenomenal field, which is associated with the perception system composed of sensation and attention, and can be forcibly updated by a perceived “fact.” A diagram of the phenomenal and perceptual fields is shown in Figure 1.
Figure 1. Phenomenal and perceptual fields.
In this illustration, the phenomenal field is precisely such a thing: it is centered on the subject, and scoped by the subject’s perceptual domain that decays with space. If there are perceived objects, this can be expressed in the following form:
(1) $f=\begin{bmatrix}N_{1}&\theta_{1}&\varphi_{1}&\gamma_{1}\\ \vdots&\vdots&\vdots&\vdots\\ N_{m}&\theta_{m}&\varphi_{m}&\gamma_{m}\\ \end{bmatrix}$
$N$ represents the meaning of the monad. This is a typical two-dimensional matrix of the form m × 4. However, it should be noted that the objects perceived in the phenomenal field are not equivalent; they have their own meaning, and the connections between them remain to be mined. Therefore, it cannot simply be characterized in a one-hot characterization using $N$.
We can easily associate word vectors. Embedding is designed to provide one-hot stronger semantic features, which can describe the features of different dimensions of words. These features are trained and learned through the positional relationships of words in a large amount of text. Therefore, a transformation is needed to expand $N$ into a specific $n$-dimensional vector that is sufficient to calculate the distance between object meanings. It can thus be unfolded in the following form after training:
(2) $f=\begin{bmatrix}N_{1}^{1}&\dots&N_{1}^{n}&\theta_{1}&\varphi_{1}&\gamma_{1}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ N_{m}^{1}&\dots&N_{m}^{n}&\theta_{m}&\varphi_{m}&\gamma_{m}\\ \end{bmatrix}$
Obviously, consciousness is a continuous process, but in a simulation, we can think of one $f$-representation as a sampling of the originally continuous consciousness at time $t$.
3.1.2. Stream of Consciousness and Internal Time-Consciousness
@@
This seems to provide an overly complex structure for every moment of experience in consciousness (i.e., as one extremely long, unbroken event), and retention is complete despite the length of the stream. As time passes, most past objects and experiences quickly lose their distinctiveness and uniqueness; ultimately, they lose any importance that deserves attention. This structure is shown in Figure 2.
Figure 2. Consciousness channel composed of retention, primal impression, and protention.
The components of retention, primal impression, and protention are all $f$-representations. In the calculation, both retention and primal impression can be obtained, but it seems that we cannot obtain a protention pointing to the future. It should be emphasized that protention is not equivalent to a primal impression at time $t$ + 1, but through retention and primal impression, consciousness can predict the next moment [37]. Just as we can naturally draw the next segment along a continuous function. Therefore, the process of obtaining protention can be represented as follows:
(3) $\begin{bmatrix}re^{0}&re^{1}&\dots&re^{t-1}&PI^{t}\end{bmatrix}\rightarrow Pro^{t},$

### v2

@@ Abstract.
Large Language Models (LLMs) still face challenges in tasks requiring understanding implicit instructions and applying common-sense knowledge. In such scenarios, LLMs may require multiple attempts to achieve human-level performance, potentially leading to inaccurate responses or inferences in practical environments, affecting their long-term consistency and behavior. This paper introduces the Internal Time-Consciousness Machine (ITCM), a computational consciousness structure to simulate the process of human consciousness. We further propose the ITCM-based Agent (ITCMA), which supports action generation and reasoning in open-world settings, and can independently complete tasks. ITCMA enhances LLMs’ ability to understand implicit instructions and apply common-sense knowledge by considering agents’ interaction and reasoning with the environment. Evaluations in the Alfworld environment show that trained ITCMA outperforms the state-of-the-art (SOTA) by 9% on the seen set. Even untrained ITCMA achieves a 96% task completion rate on the seen set, 5% higher than SOTA, indicating its superiority over traditional intelligent agents in utility and generalization. In real-world tasks with quadruped robots, the untrained ITCMA achieves an 85% task completion rate, which is close to its performance in the unseen set, demonstrating its comparable utility and universality in real-world settings.
@@ 1. Introduction
The foundation of such ability is a novel agent architecture that combines LLMs with mechanisms for synthesizing and retrieving relevant information, thus providing conditions for the output of language models. Without these mechanisms, LLMs could still output behavior for agents, but the agents might not react correctly based on past experiences, fail to make crucial inferences, or struggle to maintain long-term consistency [3].
@@
As the interaction cycles of LLM-based agents grow, two issues arise with this structure. The first issue relates to the length of the historical record. LLM-based agents process previous interactions in natural language format, appending the historical record to each subsequent input. As these records extend, they might exceed the constraints of the Transformer architectures most LLM-based agents rely on. In such cases, the system might truncate certain contents that were input into the LLM in the early stages. The second issue is the difficulty of extracting relevant memories. As agents accumulate a large volume of historical observations and action sequences, their memory burden increases continuously. This makes establishing connections between relevant topics increasingly challenging, potentially resulting in the agent’s responses being inconsistent with the ongoing context [4].
Due to the current limitations of LLMs acting as the ”brain,” recently research on LLM-based agents focuses on augmenting them with an additional layer of structures, such as LangChain [5], to optimize them. However, although these studies demonstrate the potential to better utilize LLMs through structural modifications, and some even incorporate modeling of human cognitive processes [6], the emergence of processes similar to human consciousness remains largely unexplored in existing research. This is mainly because these studies typically focus on solving specific tasks—tasks that are often specially defined and require relevant datasets to train the agents. To some extent, this sacrifices the excellent explore and exploit (EE) capabilities of generative agents, as their utility relies on the training outcomes within specific environments.
To enable generative agents to possess sufficient generalization capabilities and adapt to new environments without prior training, we believe the key lies in simulating a human-like consciousness process. This is distinct from the cognitive processes commonly discussed in the field of artificial intelligence. An agent with a similar consciousness process is not modeled as a ”rational economic agent” but rather as a ”bounded rationality” model. This means that, an agent makes decisions by constructing and testing a relatively small number of possible strategies, then selecting behaviors that sufficiently balance intuition and efficacy [35].
The most typical work here comes from cognitive science and philosophy of mind, a branch of them pays special attention to the problem of consciousness. Researchers in it discuss the problem of consciousness and human intelligence and often draw inspiration from phenomenology, especially Merleau-Ponty’s phenomenology of perception [9; 10; 11]. Interdisciplinary research in this area has revealed that intelligence relies on consciousness processes to exist. That is, intelligence is not the stimulus-response mode claimed by behaviorism (represented by reinforcement learning in AI research), nor is it an ”emergence from neuroelectricity” claimed by connectionism (represented by deep learning in artificial intelligence research). An expression of mechanical determinism is not applicable in the study of intelligent structures [12; 13].
Therefore, we believe that, in order to design a more effective generative agent architecture, it is necessary to consider a framework that simulates the human consciousness process. In this paper, we propose a computational consciousness structure for generative agents, called the Internal Time-Consciousness Machine (ITCM). It is particularly worth mentioning that the computational structure of consciousness here does not equate to generating a structure equivalent to human consciousness in a computer. Instead, it refers to a simple and abstract model, like the Conscious Turing Machine [32], which aims to algorithmically simulate the observed processes of consciousness. The basic unit of ITCM, the field, encompasses the spatial topology of perceived objects. The ITCM consists of three modules: (1) a consciousness channel module that constructs a temporal prediction process for the field to obtain a possible future change in the environment after action, (2) a memory module that aids in understanding the current environment, and (3) a driving force module that provides selection weights for the action space based on computations from the other modules. Subsequently, we designed a generative agent architecture based on ITCM, called the ITCM-based agent (ITCMA). This architecture aims to optimize the mechanism by which generative agents generate behavioral outputs based on their environment and experiences, and enhance the EE and generalization abilities of generative agents. Our main contributions are as follows:
@@
  • We propose the ITCM as a reconstruction of the underlying architecture of generative agents. This computational consciousness structure can assist generative agents in handling complex tasks with greater flexibility and intelligence while enhancing their interpretability, making their behaviors easier to understand and predict.
@@
  • Technically, based on the proposed ITCM structure, we introduce an ITCMA and validate its effectiveness in both of the real life scenario environments described by text, as well as in real-world robotic environments. Based on the ITCM structure, the ITCMA demonstrated strong EE and generalization abilities in experimental environments, achieving high utility even without prior training in these environments.
@@ 2.1. LLMs as agents
The emergence of LLMs has brought new avenues for intelligent agents. In recent years, there has been a proliferation of LLM-based agent architectures [14; 15; 4]. These architectures primarily focus on two aspects: planning and tools [16].
@@
However, as previously mentioned, current research focuses on improving the training effectiveness of generative agents on specific tasks through enhancements to planning and tool modules. This sacrifices some of the superiority of generative agent structures in EE performance. In this paper, we aim to enhance the EE and generalization abilities of generative agents by proposing the Internal Time-Consciousness Machine (ITCM), a computational consciousness structure adapted for generative agents. ITCM simulates human consciousness processes rather than cognitive processes. This is achieved by reconstructing the underlying logic of generative agents, thereby improving their EE and generalization abilities.
@@ 2.2.1. Integrated Information Theory
The central claim of integrated information theory (IIT) is that a physical system is conscious when and only when it is the maximum of the integrated information $\phi$ [27; 28]. As defined by the IIT, integrated information can be roughly described as a measure of the extent to which a system is causally bound to its own past and future states and the dependence of these constraints on the causal interconnectedness between system components. If a system has more integrated information than any overlapping system (e.g., a smaller system that is a part of it or any larger system containing it), then this system has the maximum comprehensive information. Consciousness needs the largest $\phi$, not just a non-zero value of $\phi$ (this is called the exclusion theorem). The IIT also claims that consciousness is identical to the maximum $\phi$.
The IIT has been met with some philosophical objections [29]. One of these objections is that consciousness intuitively looks like an intrinsic property, but maximal $\phi$ is an extrinsic property. Therefore, if this intuition is correct, then consciousness cannot be the same as maximal $\phi$ [30].
@@ 2.2.2. Consciousness Turing Machine
According to Blum [32], CTM is not meant to model the brain or imply neural correlates of consciousness. It is a simple abstract model of consciousness aimed at understanding consciousness and its related phenomena. Therefore, the definition of CTM adopts a theoretical computer science perspective, forming a conscious Global Workspace Theory (GWT) mathematically.
@@
Therefore,the basis of consciousness in the CTM uses Baar’s theater of consciousness hypothesis, which likens consciousness to the performance of dramatic actors on a stage of working memory. Their performance occurs under the observation of a group of spectators sitting in the dark (unconscious processors). The stage of the GWT is represented by short-term memory (STM) that contains the content of the CTM consciousness at any moment. The spectators are represented by powerful processors. Each processor has its own expertise, and these processors make up the CTM’s long-term memory (LTM). These processors make predictions and receive feedback from the CTM’s world. The learning algorithms inside each processor improve the processor’s behavior based on this feedback.
When all LTM processors receive these contents via this broadcast, CTM becomes consciously aware of these contents. In CTM, the definition of consciousness roughly aligns with what cognitive neuroscientists refer to as ”attention.” The feeling of consciousness in CTM corresponds roughly to what cognitive neuroscientists refer to as ”consciousness” or ”subjective consciousness.”
3. The Internal Time-Consciousness Machine based Agent (ITCMA)
@@
It is undeniable that the IIT and CTM suggest ways to build algorithms based on different aspects of consciousness research. However, the question remains as to whether there is truly a computation-based theory of consciousness, even though there are many similarities between the brain and the structure of the digital computer. Thagard [33] summarizes seven challenges to the study of representation computation, and the asymmetry between cognition and computation occupies the first position in this list of challenges. In Appendix A, We discuss in detail the necessity of focusing on constructing a pseudo-consciousness model corresponding to intelligence by turning our attention to first-person phenomenology. However, at present, we need to first introduce the overall framework of the internal time-consciousness machine based agent (ITCMA) as a generative agent structure. As for how the internal time-consciousness machine (ITCM), which is the computational consciousness structure behind ITCMA, is proposed in accordance with phenomenological thinking, we will discuss it in Appendix B. In the following part of this section, we will mainly discuss the model form of ITCM.
3.1. The overall structure of ITCMA
Park et al. [3] proposed the original generative agent structure in which all content is recorded and inferred in natural language, enabling the structure to utilize an LLM. Influenced by this, Xi et al. [4] proposed a general conceptual framework for generative agents that includes three key components: the brain, perception, and action.
The structure of the ITCMA is similar in that perception serves as input from the environment, and action serves as output to the environment. After receiving action, the environment forces the agent to form a perception as a re-action to it. Reaction is the carrier of this action—the brain is not equivalent to the LLM here but is replaced by the complete ITCMA, and the LLM is only used as a tool. The structure is shown in Figure 1.
Figure 1. Conceptual framework of ITCMA.
Figure 1 is based on the specific ITCMA structure, which can be further refined. Based on the foundational structure of the brain, perception, and action, we first define the basic perceptual unit of ITCMA, the field, which includes the spatial topology of perceived objects. Following this, we define three modules for ITCMA:
@@ 3.1. The overall structure of ITCMA
    The consciousness channel module, which constructs the temporal prediction process for the field.
@@
    The memory module, which aids in understanding the current environment.
@@
    The driving force module, which provides selection weights for the action space based on computations from the other modules.
The interaction between these three modules follows the following process: after receiving environmental information, the consciousness channel module first changes some of its own content (retention and primal impression), and inputs this part to the memory module to obtain the activated memory. This memory will be used to generate the remaining part of the consciousness channel module (Protention). After that, the contents of these two modules will be input to the driving force module to calculate the weight of the action space, and then the LLM will make a decision on the action.
In more detailed expressions, at time $t$, the current perception provided by environment (primal impression) $PI^{t}$ and the list retention $Re^{t}$ made up of perceptions from previous moments are combined. The ITCMA uses this to extract the activated memory $AM$ and its subsequent content from long-term memory. These three together form consciousness channel $C^{t}$, at time $t$ after receiving $C^{t-1}$ of time $t-1$. Through $C^{t}$, ITCMA can use a time series forecasting model (TSFM) to calculate for each action in the action space, in order to obtain the possible changes in the environment (Protention) $Pro^{t}$ after they are executed. It will then be used to calculate the selection weight driving force $d^{t}$ for each action. The above process is ultimately provided to the LLM in natural language to select an action to execute. The entire process is shown in Figure 2.
Figure 2. Detailed structure of the ITCMA.
Even though LLM is not used as a direct brain like in other generative agent structures, its role in ITCMA is also very central. Its role is to understand the information provided by the perception module and, based on this information, combine its own understanding of the world to output an action that the agent will execute. Therefore, the working process of ITCMA can be compared to a few-shot CoT that mimics the human consciousness thinking process. The ITCM structure can give LLM a regular human-like thinking logic through an explore and exploit (EE) process and a context model centered on the extracted more suitable environmental memory, thereby establishing a new rule system beyond LLM.
In addition, due to the ITCM constructing an agent that simulates consciousness without experience, using a visual language model (VLM) to aid in modeling is an option here. In complex task environments, the information captured by vision is not only the spatial topology of an object but also includes a holistic understanding of it. Using the VLM here to characterize the visual image in natural language form and add it to the ITCMA workflow can somewhat bypass its early process of understanding of spatial relationships.
However, due to the enhancement of the VLM’s unique visual ability, its language generation ability is slightly insufficient compared to an LLM of the same scale. Therefore, it is only used as a descriptive tool here and cannot replace the original LLM used to generate actions. In an evaluation based on real-world quadrupedal robot environments, we will use the VLM to accomplish an understanding of three-dimensional spatial information.
In the next section, we will introduce the details and mechanisms of each module. Appendix C includes the specific implementation details and examples of the ITCMA.
3.2. Modules in ITCM
3.2.1. Field
A ”field” is a functional and exclusive activity location of conscious experience, as identified by Husserl [36]. According to Merleau-Ponty [12], the perceptual subject is always in the midst of other objects; it is always a part of the field. Without a mechanism such as the perception field, our perception would not be continuous at all. Therefore, let us imagine a spherical coordinate system: the perceptual subject (that is, the agent) is always in the middle of the sphere and facing forward, and the object that can be perceived is described by the set of spherical coordinates ($\theta$, $\varphi$, $\gamma$). When we obtain these sets of represented objects, the spherical coordinate system that can be transformed by the agent’s actions becomes the agent’s phenomenal field.
A field is therefore decomposed into two parts: a perceptual field and a phenomenal field. Among them, the phenomenal field is a spherical coordinate system centered on the subject, which contains many objects, and it can make some predictive changes to these objects according to experience; the perceptual field is a part of the phenomenal field, which is includes those objects that are directly perceived. The objects contained by it can be forcibly updated by a perceived ”fact.” For example, during a speech, the speakers can directly perceive the audience’s reactions and also consider, through imagination, the possible discussions among audience members that they cannot directly hear. A diagram of the phenomenal and perceptual fields is shown in Figure 3.
Figure 3. Phenomenal and perceptual fields. The white area represents the phenomenal field, within which object 3 has not been perceived by the agent. The gray area represents the perceptual field, within which objects 1 and 2 are directly perceived by the agent.
In this illustration, the phenomenal field is precisely such a spherical coordinate: it is centered on the agent, and scoped by the agent’s perceptual domain that decays with space. If there are perceived objects, they contain two attributes: an attribute $N$ representing their own content and an attribute $pos\leftarrow\begin{bmatrix}\theta&\varphi&\gamma\end{bmatrix}$ representing their own position. For $N$, we can easily associate word vectors. A transformation is needed to expand $N$ into a specific -dimensional vector that is sufficient to calculate the distance between object meanings. Therefore, for a field containing $m$ objects, it can be unfolded in the following form after training:
(1) $f=\begin{bmatrix}N_{1}&\theta_{1}&\varphi_{1}&\gamma_{1}\\ \vdots&\vdots&\vdots&\vdots\\ N_{m}&\theta_{m}&\varphi_{m}&\gamma_{m}\\ \end{bmatrix}=\begin{bmatrix}N_{1}^{1}&\dots&N_{1}^{n}&\theta_{1}&\varphi_{1}&\gamma_{1}\\ \vdots&\ddots&\vdots&\vdots&\vdots&\vdots\\ N_{m}^{1}&\dots&N_{m}^{n}&\theta_{m}&\varphi_{m}&\gamma_{m}\\ \end{bmatrix}$
3.2.2. Consciousness Channel Module
The field provides a topological modeling of the perceived objects. However, consciousness is a continuous process. We can think of one $f$-representation as a sampling of the originally continuous consciousness at time $t$. By establishing a continuous consciousness stream channel through multiple frames, we can construct a temporal structure that will be used for future predictions.
According to Husserl [36], the basic unit of temporality is a temporal field that comprises all three temporal modes of present, past, and future. The modeling of the past, we refer to as (1) Retention. It consists of several temporally continuous fields. It should be emphasized that retention should be distinguished from recall. There is a significant difference between a person’s impression of something that just happened (retention) and a person’s memory of a past event. A retentionis directly connected to a (2) Primal impression. Primal impression is a single field that represents the raw content of information perceived at the present moment. By combining the temporal elements of retention and primal impression, we can predict how the field will change in the next moment, thus obtaining a direction towards future information. The possible change state of the field at the next moment after the execution of the action is called (3) Protention. It should be emphasized that protention is not equivalent to a primal impression at time $t+1$, but through retention and primal impression, consciousness can predict the next moment [39]. Just as we can naturally draw the next segment along a continuous function. Therefore, the process of obtaining protention can be represented as follows:
(2) $\begin{bmatrix}Pro^{t}\leftarrow re^{0}&re^{1}&\dots&re^{t-1}&PI^{t}\end{bmatrix},$